OpenStudy (anonymous):

Use the quadratic formula to find the solution to the quadratic equation below. x^2 - 5x + 2

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] Apply the quadratic formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (unklerhaukus):

so in your case \[a=1,\qquad b=-5,\qquad c=2.\]

OpenStudy (unklerhaukus):

\[x_{1,2}=\frac{-(-5)\pm\sqrt{(-5)^2-4(1)(2)}}{2(1)}\\\qquad=\frac{5\pm\sqrt{25-8}}{2}\\\qquad=\]

OpenStudy (anonymous):

these are my answers to choose from A. 5±√33 ________ 2 B. 5±√-8 ________ 2 C. 5±√17 _______ 2 D. -2±√21 ________ 2 E. -5±√3 ________ 2 F. 5±√-8 ________ 2

OpenStudy (unklerhaukus):

25-8=

OpenStudy (anonymous):

17 so my answer would be C?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (whpalmer4):

Look, @UnkleRhaukus did the problem for you, you could at least select the right answer from his work :-)

OpenStudy (whpalmer4):

Do any of the other choices look at all like they might be possible, given what he's done? It really is important to be able to make a calculation and have some confidence in the result. D and E can be stricken from consideration due to having something other than \(5 \pm\) at the start of the numerator. \(\sqrt{25-8}\) simplifies to what? Hopefully you agree that \(\sqrt{-8}\) isn't a possibility, so that rules out B and F (do you really have the same answer choice twice?) 25+8 = 33, so you should look carefully at your work and make sure you (or your helper) haven't made a simple mistake with a sign somewhere. Then when you have done so, make your best choice. As a dear old math teacher of mine used to say when someone answered a question like that, "is that an answer, or a prayer?" :-)

OpenStudy (whpalmer4):

You'll make mistakes. We all do. That's okay. When you do make a mistake, go back and figure out why you made the mistake. Didn't read the problem statement or answer choices quite carefully enough? Next time, slow down and take the time to do so. Sloppy handling of negative signs? Check everything you do with negative signs, maybe underline terms that have them as you go along as a bigger visual signal.